# Dark Matter Physics in General NMSSM

### v1

@@ 3.2 Global features
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: Two-dimensional distributions of the marginal posterior PDF for the likelihood $\mathcal{L}$ in Eq. 3.3, projected onto the $|m_{\tilde{\chi}_{1}^{0}}|-\sigma^{\rm SI}_{\tilde{\chi}_{1}^{0}-p}$ and $|m_{\tilde{\chi}_{1}^{0}}|-\sigma^{\rm SI}_{\tilde{\chi}_{1}^{0}-n}$ planes. The left panels show the results of the $h_{1}$ scenario, while the right panels show those of the $h_{2}$ scenario. The solid and dashed lines surround the $1\sigma$ and $2\sigma$ credible regions, respectively, and the samples falling within these regions contributed $65.3\%$ and $95.4\%$ of the total Bayesian evidence [89]. All the considered samples were consistent with the LHC constraint.
@@ 3.3 Three annihilation channels
  • The s-wave nature of the annihilation $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$ with $h_{s}\to b\bar{b}$ and $A_{s}\to b\bar{b}$ also allows for indirect DM search experiments to establish a lower bound on $|m_{\tilde{\chi}_{1}^{0}}|$ (approximately $58~{}{\rm GeV}$). This bound is substantially less stringent than that of the direct annihilation $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to b\bar{b}$, as the power produced through the former channel is suppressed by its relatively heavy DM mass [138, 139]. In contrast, no experimental limits apply to p-wave annihilations $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}h_{s}$ and $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to A_{s}A_{s}$, permitting potentially much lighter DM candidates down to ten GeV.
@@ 4 Summary
Second, we examined the characteristics of Singlino-dominated DM and identified significant distinctions from the $\mathbb{Z}_{3}$-NMSSM. In the latter, DM properties were determined by four parameters: $\lambda$, $\tan\beta$, $\mu_{\rm eff}$, and $m_{\tilde{\chi}_{1}^{0}}\simeq 2\kappa\mu_{\rm eff}/\lambda$. For a Singlino-dominated neutralino to remain the LSP, $|\kappa|$ had to be smaller than $\lambda/2$. Furthermore, the LZ constraint on $\lambda$ necessitated $|\kappa|$ to be very small. As a result, the Singlino DM in the $\mathbb{Z}$-NMSSM needed to co-annihilate with electroweakinos to match the observed relic abundance. However, such dynamics differed in the GNMSSM due to the introduction of $\mathbb{Z}_{3}$-violating terms where the DM’s properties were described through five independent parameters: $\lambda$, $\kappa$, $\tan\beta$, $\mu_{tot}$, and $m_{\tilde{\chi}_{1}^{0}}$. Specifically, allowing for decorrelation of $m_{\tilde{\chi}_{1}^{0}}$ from $\mu_{tot}$ enables $|\kappa|$ to exceed $\lambda$ when predicting Singlino-dominant DM. This structural difference significantly impacted the potential creation of a secluded DM sector within singlet particles [45]. The key characteristics of the secluded sector in our analysis included the following:
@@
  • The s-wave nature of $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$ resulted in a lower limit on $m_{\tilde{\chi}_{1}^{0}}$ at approximately $60~{}{\rm GeV}$ from indirect DM search experiments, whereas in other channels, $m_{\tilde{\chi}_{1}^{0}}$ could decrease to as low as ten GeV.

### v2

@@ 3.2 Global features
Figure 4: Two-dimensional distributions of the marginal posterior PDF for the likelihood $\mathcal{L}$ in Eq. 3.3, projected onto the $|m_{\tilde{\chi}_{1}^{0}}|-\sigma^{\rm SI}_{\tilde{\chi}_{1}^{0}-p}$ and $|m_{\tilde{\chi}_{1}^{0}}|-\sigma^{\rm SD}_{\tilde{\chi}_{1}^{0}-n}$ planes. The left panels show the results of the $h_{1}$ scenario, while the right panels show those of the $h_{2}$ scenario. The solid and dashed lines surround the $1\sigma$ and $2\sigma$ credible regions, respectively, and the samples falling within these regions contributed $65.3\%$ and $95.4\%$ of the total Bayesian evidence [95]. All the considered samples were consistent with the LHC constraint.
@@ 3.3 Three annihilation channels
  • The s-wave nature of the annihilation $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$ with $h_{s}\to b\bar{b}$ and $A_{s}\to b\bar{b}$ also allows for indirect DM search experiments to establish a lower bound on $|m_{\tilde{\chi}_{1}^{0}}|$ (approximately $58~{}{\rm GeV}$). This bound is substantially less stringent than that of the direct annihilation $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to b\bar{b}$, as the power produced through the former channel is suppressed by its relatively heavy DM mass [144, 145]. In contrast, no experimental limits apply to p-wave annihilations $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}h_{s}$ and $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to A_{s}A_{s}$, permitting potentially much lighter DM candidates down to 10 GeV.
@@ 4 Summary
Second, we examined the characteristics of Singlino-dominated DM and identified significant distinctions from the $\mathbb{Z}_{3}$-NMSSM. In the latter, DM properties were determined by four parameters: $\lambda$, $\tan\beta$, $\mu_{\rm eff}$, and $m_{\tilde{\chi}_{1}^{0}}\simeq 2\kappa\mu_{\rm eff}/\lambda$. For a Singlino-dominated neutralino to remain the LSP, $|\kappa|$ had to be smaller than $\lambda/2$. Furthermore, the LZ constraint on $\lambda$ necessitated $|\kappa|$ to be very small. As a result, the Singlino DM in the $\mathbb{Z}_{3}$-NMSSM needed to co-annihilate with electroweakinos to match the observed relic abundance. However, such dynamics differed in the GNMSSM due to the introduction of $\mathbb{Z}_{3}$-violating terms where the DM’s properties were described through five independent parameters: $\lambda$, $\kappa$, $\tan\beta$, $\mu_{tot}$, and $m_{\tilde{\chi}_{1}^{0}}$. Specifically, allowing for decorrelation of $m_{\tilde{\chi}_{1}^{0}}$ from $\mu_{tot}$ enables $|\kappa|$ to exceed $\lambda$ when predicting Singlino-dominant DM. This structural difference significantly impacted the potential creation of a secluded DM sector within singlet particles [51]. The key characteristics of the secluded sector in our analysis included the following:
@@
  • The s-wave nature of $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$ resulted in a lower limit on $m_{\tilde{\chi}_{1}^{0}}$ at approximately $60~{}{\rm GeV}$ from indirect DM search experiments, whereas in other channels, $m_{\tilde{\chi}_{1}^{0}}$ could decrease to as low as 10 GeV.